OpenStudy (anonymous):

Write a direct variation equation that relates x to y. Then solve. Show both the equation and the solution. If y = 8 when x = 4, find y when x = 5.

OpenStudy (anonymous):

if y=8 and x=4; we have to find y when x=5 the equation is, if y=z*4 8=z*4 z=8/4 z=2 similarly y=z*x y=2*5 (put the value of z) y=10 Correct?

Parth (parthkohli):

Yes, you made me erase the whole answer I typed.. correct ;p

OpenStudy (savvy):

lol...